OpenStudy (anonymous):

if a throw rug 1m 2.5 m costs 54 how much does it cost per square meter

OpenStudy (anonymous):

We have to get the area first A=bh A=(1)(2.5) A=2.5 m^2 2.5m^2= 54 Divide both sides by 2.5 1m^2=21.6